Solve the Partial differential equation

x² d²z/dx² - y² d²z/dy² + x dz/dx - y dz/dy = ln(x)
1
Expert's answer
2019-08-26T09:18:42-0400

Let "x=e^t", then

"ln(x)=t"

Using chain rule:

"\frac{\partial z}{\partial t}=\frac{\partial z}{\partial x}\frac{dx}{dt}=\frac{\partial z}{\partial x}e^t"

"\frac{\partial z}{\partial x}=\frac{\partial z}{\partial t}e^{-t}"

"\frac{\partial ^2z}{\partial t^2}=(\frac{\partial z}{\partial x}e^t)'_t=\frac{\partial ^2z}{\partial x^2}\frac{dx}{dt}e^t+\frac{\partial z}{\partial x}e^t=\\
\frac{\partial ^2z}{\partial x^2}e^{2t}+\frac{\partial z}{\partial t}"

"\frac{\partial ^2z}{d\partial ^2}=(\frac{\partial ^2z}{\partial t^2}-\frac{\partial z}{\partial t})e^{-2t}"


"x^2\frac{\partial ^2z}{\partial x^2}+x\frac{\partial z}{\partial x}=e^{2t}(\frac{\partial ^2z}{\partial t^2}-\frac{\partial z}{\partial t})e^{-2t}+e^t\frac{\partial z}{\partial t}e^{-t}=\\
\frac{\partial ^2z}{\partial t^2}-\frac{\partial z}{\partial t}+\frac{\partial z}{\partial t}=\frac{\partial ^2z}{\partial t^2}"

Similarly after substitution "y=e^s" we have

"y^2\frac{\partial ^2z}{d\partial ^2}+y\frac{\partial z}{\partial y}=\frac{\partial ^2z}{\partial s^2}"

And equation becomes:

"\frac{\partial ^2z}{\partial t^2}-\frac{\partial ^2z}{\partial s^2}=t"

Now introduce new variables "\xi=t+s,\ \eta=t-s"

"z=z(\xi,\eta)"

Using chain rule again:

"\frac{\partial z}{\partial t}=\frac{\partial z}{\partial \xi}\frac{\partial \xi}{\partial t}+\frac{\partial z}{\partial \eta}\frac{\partial \eta}{\partial t}=
\frac{\partial z}{\partial \xi}+\frac{\partial z}{\partial \eta}"

"\frac{\partial z}{\partial s}=\frac{\partial z}{\partial \xi}\frac{\partial \xi}{\partial s}+\frac{\partial z}{\partial \eta}\frac{\partial \eta}{\partial s}=
\frac{\partial z}{\partial \xi}-\frac{\partial z}{\partial \eta}"

"\frac{\partial ^2z}{\partial t^2}=\frac{\partial ^2z}{\partial \xi^2}\frac{\partial \xi}{\partial t}+\frac{\partial ^2z}{\partial \xi \partial \eta}\frac{\partial \eta}{\partial t}+\frac{\partial ^2z}{\partial \xi \partial \eta}\frac{\partial \xi}{\partial t}+\frac{\partial ^2z}{\partial \eta^2}\frac{\partial \eta}{\partial t}="

"\frac{\partial ^2z}{\partial \xi^2}+2\frac{\partial ^2z}{\partial \xi \partial \eta}+\frac{\partial ^2z}{\partial \eta^2}\frac{\partial \eta}{\partial t}"

"\frac{\partial ^2z}{\partial s^2}=\frac{\partial ^2z}{\partial \xi^2}\frac{\partial \xi}{\partial s}+\frac{\partial ^2z}{\partial \xi \partial \eta}\frac{\partial \eta}{\partial s}-(\frac{\partial ^2z}{\partial \xi \partial \eta}\frac{\partial \xi}{\partial s}+\frac{\partial ^2z}{\partial \eta^2}\frac{\partial \eta}{\partial s})="

"\frac{\partial ^2z}{\partial \xi^2}-\frac{\partial ^2z}{\partial \xi \partial \eta}-(\frac{\partial ^2z}{\partial \xi \partial \eta}-\frac{\partial ^2z}{\partial \eta^2})=
\frac{\partial ^2z}{\partial \xi^2}-2\frac{\partial ^2z}{\partial \xi \partial \eta}+\frac{\partial ^2z}{\partial \eta^2})"

"\frac{\partial ^2z}{\partial t^2}-\frac{\partial ^2z}{\partial s^2}=4\frac{\partial ^2z}{\partial \xi \partial \eta}"

"t=(t+s)/2+(t-s)/2=\frac{\xi+\eta}{2}"

"4\frac{\partial ^2z}{\partial \xi \partial \eta}=\frac{\xi+\eta}{2}"

"\frac{\partial ^2z}{\partial \xi \partial \eta}=\frac{\xi+\eta}{8}"

Integrating:

"\frac {\partial z}{\partial \eta}=\int \frac{\partial ^2z}{\partial \xi \partial \eta}\partial \xi=\int \frac{\xi+\eta}{8}d\xi=\xi^2/16+\eta\xi/8+g(\eta)"

"z=\int (\xi^2/16+\eta\xi/8+g(\eta))d\eta=\\
\xi^2\eta/16+\eta^2\xi/16+G(\eta)+F(\xi)",

where F and G are arbitrary twice differentiable functions,

and G is antiderivative of g

Now use reverse substitutions "t=ln(x),\ s=ln(y)"

we have:

"\xi=t+s=ln(x)+ln(y)=ln(xy),\\
 \eta=t-s=ln(x)-ln(y)=ln(x/y)"

"z=\frac{ln(xy)ln(x/y)}{16}(ln(xy)+ln(x/y))+G(ln(x/y)+F(ln(xy))=\\
\frac{ln(xy)ln(x/y)}{8}(ln(x))+G(ln(x/y)+F(ln(xy))"

Answer: "z=\frac{ln(xy)ln(x/y)}{8}(ln(x))+G(ln(x/y)+F(ln(xy))", where F and G are arbitrary twice differentiable functions.